OpenStudy (anonymous):

Will fan and give medal :) Which conjunction or disjunction is equivalent to the open sentence |a – 9| ≤ 4. A. a – 9 ≥ 4 and a – 9 ≤ –4 B. a – 9 ≤ 4 or a – 9 ≥ –4 C. a – 9 ≤ 4 and a – 9 ≤ –4 D. a – 9 ≤ 4 and a – 9 ≥ –4

OpenStudy (jdoe0001):

recall that \(\bf |something| \le whatever\implies \begin{cases} +something \le whatever\\ \quad \\ \bf -something \le whatever \end{cases}\)

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

@amoodarya

OpenStudy (anonymous):

I still don't understand...

OpenStudy (anonymous):

@KirbyLegs

OpenStudy (jdoe0001):

when multiplying for a negative value, in inequalities, you'd have to "flip" the inequality sign as well, so \(\bf |something| \le whatever\\ \quad \\\implies \begin{cases} +something \le whatever\\ \quad \\ \bf -something \le whatever\implies something \ge -whatever \end{cases}\) plug in your values given in \(\bf |a-9|\le 4\) and see what you get

OpenStudy (anonymous):

okay thanks :)

OpenStudy (jdoe0001):

yw